OpenStudy (anonymous):

Karen makes $5 per hour baby-sitting and $12 per hour giving music lessons. One weekend, she worked a total of 18 hours and made $139. How many hours did she spend baby-sitting?

OpenStudy (anonymous):

This shall be funn Let x = number of hours babysitting Let y = number of hours giving lessons From the information, you can write two equations x + y = 18 5x + 12y = 139 Solve for x in the first equation : x + y = 18 x = 18 - y Put that value of x in the second equation and solve for y : 5x + 12y = 139 5(18 - y) + 12y = 139 90 - 5y + 12y = 139 90 + 7y = 139 7y = 49 . . . . . . . . . . . . . . [ subtracted 90 from both sides ] y = 7 . . . . . . . . . . . . . . . .[ divided both sides by 7 ] Now go back to the very first equation and put that value of y in and solve for x : x + y = 18 x + 7 = 18 x = 11 She spent 11 hours babysitting. ------ Check: (11 hours x $5) + (7 hours x $12) = ($55 + $84) = $139

OpenStudy (anonymous):

Did you understand that?

OpenStudy (anonymous):

Ummm , kinda / not really :D

OpenStudy (anonymous):

lols haha is there anything you need me to clarify?

OpenStudy (anonymous):

This part :D Put that value of x in the second equation and solve for y : 5x + 12y = 139 5(18 - y) + 12y = 139 90 - 5y + 12y = 139 90 + 7y = 139 7y = 49 . . . . . . . . . . . . . . [ subtracted 90 from both sides ] y = 7 . . . . . . . . . . . . . . . .[ divided both sides by 7 ]